# Hydrodynamic bend instability of motile particles on a substrate

### v2

@@ Abstract
[…]
@@
where, $f$ is the friction coefficient. $\alpha{\bm{p}}$ is the only active term characterizing the self propulsion in the system that describes the migration of active particles along their polarity direction ${\bm{p}}$, and the coefficient $\alpha$ sets the self-propulsion strength. In the absence of any active force, the term $\beta{\bm{h}}$ is required to fulfill the detailed balance condition compensating for the $\beta{\bm{v}}$ term in Eq. (1) [50]. The stress tensor, $\sigma_{ij}$, encompasses both the viscous stress, $\sigma_{ij}^{\text{viscous}}=2\eta E_{ij}$ with $\eta$ being the viscosity coefficient, and the elastic stress $\sigma_{ij}^{\text{elastic}}=-P\delta_{ij}+\frac{1-\lambda}{2}p_{i}h_{j}-\frac{1+\lambda}{2}p_{j}h_{i}+\frac{\lambda}{2}p_{k}h_{k}\delta_{ij}$, with $P$ being the isotropic pressure [56]. We omit the coupling to the concentration field, as it does not qualitatively alter the characteristic states observed, nor does it affect the instability, see Appendix Hydrodynamic bend instability of motile particles on a substrate.
@@
Linear stability analysis:- To rigorously characterize the observed hydrodynamic bend instability driven solely by self-propulsion force, we examine the stability of the minimal model that couples the following dynamic variables: the two components of the polarization, $p_{x}$ and $p_{y}$, and the vorticity, $\Omega$. The vorticity is defined as $\Omega=\partial_{x}v_{y}-\partial_{y}v_{x}$ and alongside the steam function $\psi$ given by $v_{x}=\partial_{y}\psi$ and $v_{y}=-\partial_{x}\psi$ ensures the incompressibility condition. To isolate the essential mechanism underlying the instability, we neglect co-rotational terms that couple polarity dynamics to velocity gradients, as well as advective terms in the polarity and velocity equations. This simplification allows us to focus on the interplay between self-propulsion and flow alignment with the minimal ingredients required to trigger the instability. In the Supplementary Information [57], we show that including the full terms does not suppress the instability, underscoring the robustness of the underlying mechanism. Under these conditions, one can then write the Navier-Stokes equation (Eq. 2) in terms of vorticity
@@
This relation describes the evolution of the perturbation through a non-linear relation between the wave-number $q$ and the growth rate $\tilde{\omega}$. By differentiating Eq. (8) with respect to the wavenumber, we find that the growth rate reaches its maximum at the smallest wavenumber, $q=0$, similar to the hydrodynamic instabilities generated by dipolar flows in both polar [56] and nematic [58, 59] active systems. The fastest-growing mode corresponds to the wavenumber $q=\frac{2\pi}{L}$, specifically the longitudinal mode, with $n=1$ and $m=0$. In Fig. 3 we display the growth rate curves, numerically solved from Eq. (8), for a constant self-propulsion coefficient and varying $\beta$. The critical self-propulsion $\alpha_{c}$ that separates the stable ($\tilde{\omega}<0$) from the unstable ($\tilde{\omega}>0$) states can be found by examining when $\tilde{\omega}$ crosses zero. This leads to the following relation
${\color[rgb]{0,0,0}\alpha_{c}=Kq^{2}\left(\frac{f}{\gamma}\frac{1}{\beta}+\beta\right),}$ (9)
@@
Mechanism:- The relation that determines the systems stability, given by Eq. (9), reveals the three key components driving this instability: (i) the flow-aligning parameter $\beta$, (ii) the self-propulsion strength $\alpha$, and (iii) the Frank elastic constant $K$. Here, self-propulsion aligns the velocity with the polarization, and conversely, $\beta$ aligns the polarization with the velocity (see schematic in Fig. 4a). When spatial variations in polarity occur, a competition arises between these aligning tendencies and the Frank elastic energy, which penalizes gradients in the polarization field. To illustrate this, consider an initially aligned velocity field with a polarity exhibiting a gradient along the $x$-axis (Fig. 4c). The self-propulsion and flow-alignment terms act to align velocity and polarity locally (Fig. 4d). However, this alignment frustrates the Frank elasticity, which favors a uniform polarization direction to minimize distortions. This competition creates a feedback loop: local misalignments lead to velocity changes that further distort the polarity field, amplifying fluctuations rather than suppressing them. Importantly, the instability arises because the flow alignment $\beta$ couples the polarity and velocity fields dynamically, and the absence of instantaneous force transmission allows these distortions to grow.
Discussion:- Our results demonstrate that a suspension of active self-propelled particles that are in contact with a substrate, the self-propulsive force alone is enough - in the absence of any dipolar active stress - to induce bend instabilities. This is interesting, since the hydrodynamics bend instability of a suspension of active particles is typically associated with active dipolar stresses.
@@
We note an important distinction between our results and those reported in Ref. [50], which reports that long-range hydrodynamic interactions suppress fluctuations in the ordered state, even when noise and nonlinearities are included. While this appears at odds with the bend instability we report, the discrepancy arises from fundamentally different dynamical assumptions. Ref. [50] effectively assumes an adiabatic dynamics in the quasi-static force balance regime in which polarity relaxes much more slowly than mechanical variables, such that the polarity field is slaved to instantaneous force balance. This assumption suppresses dynamical feedbacks between polarity evolution, self-propulsion, and flow. In contrast, our model retains the full coupled dynamics of polarity, flow, and self-propulsion, without assuming time-scale separation. This more general framework, employed in several works on active nematics and polar fluids [35, 60, 56, 59, 61, 62], reveals an instability that emerges from the dynamical competition between alignment to flow and self-propulsion.
@@
Our findings are potentially applicable to hydrodynamic instabilities and the early stages of collective motion in systems of self-propelled particles, such as motility assays of polar filaments [13], migratory cells in monolayers [51, 10, 63], motile bacteria in two-dimensional setups [64, 65, 66], and active colloids [12, 38, 67] that move on substrates and tend to align with the local flow direction. In such systems, flow-alignment interactions are not merely a modeling assumption but arise naturally due to the presence of a physical substrate, which mediates frictional coupling and reinforces alignment between polarity and flow [68]. While prior work has often interpreted observed instabilities primarily through the lens of active stress–driven mechanisms, our analysis reveals a complementary and robust route to bend instabilities emerging from the interplay of self-propulsion and flow alignment. It is important to emphasize that this mechanism is distinct from the classic active nematic instability, which arises from gradients of dipolar active stress that generate destabilizing flows. These two routes to instability, one rooted in self-propulsion and flow alignment, the other in dipolar stress and flow gradient-alignment feedback, may coexist or compete, depending on system parameters. Distinguishing between them requires careful analysis of polarity dynamics and their relation to local flow versus strain rate fields. In this context, individual-based hydrodynamic models such as squirmers [69], which allow for independent tuning of self-propulsion and dipolar activity [70], offer a promising platform for systematically probing the interplay between these mechanisms. Likewise, direct measurements of active forces in experiments could help elucidate the relative contributions of polar and dipolar activity to the observed dynamics.
@@ Appendix B Concentration effect on the linear stability
[…]

### v3

@@ Abstract
We use general dynamical equations for the polarization field, ${\bm{p}}({\bm{r}},t)$, of active units suspended in a fluid with the total velocity field of the particles and the fluid being ${\bm{v}}({\bm{r}},t)$, where $\bm{r}$} is a two-dimensional position vector. The joint density, $\rho$ of the particles and the fluid is conserved, i.e., $\dot{\rho}=0$ implying $\partial_{i}v_{i}=0$. In the absence of activity and fluid flow, the equilibrium relaxation derives from a Landau–de Gennes free energy given as, $\mathcal{F}=\int dA[C(1-\lvert p\rvert^{2})^{2}+\frac{K}{2}|\nabla{\bm{p}}|^{2}]$, where $C$ sets the energy scale for Landau-de Gennes free energy [52]and $K$ is the Frank elastic constant within the single elastic constant approximation [53]. The equation for polarization, $\bm{p}$} is,
@@
where, $f$ is the friction coefficient. $\alpha{\bm{p}}$ is the only active term characterizing the self propulsion in the system that describes the migration of active particles along their polarity direction $\bm{p}$}, and the coefficient $\alpha$ sets the self-propulsion strength. In the absence of any active force, the term $\beta{\bm{h}}$ is required to fulfill the detailed balance condition compensating for the $\beta{\bm{v}}$ term in Eq. (1) [50]. The stress tensor, $\sigma_{ij}$, encompasses both the viscous stress, $\sigma_{ij}^{\text{viscous}}=2\eta E_{ij}$ with $\eta$ being the viscosity coefficient, and the elastic stress $\sigma_{ij}^{\text{elastic}}=-P\delta_{ij}+\frac{1-\lambda}{2}p_{i}h_{j}-\frac{1+\lambda}{2}p_{j}h_{i}+\frac{\lambda}{2}p_{k}h_{k}\delta_{ij}$, with $P$ being the isotropic pressure [56]. We omit the coupling to the concentration field, as it does not qualitatively alter the characteristic states observed, nor does it affect the instability, see Appendix B.
@@
Linear stability analysis:- To rigorously characterize the observed hydrodynamic bend instability driven solely by self-propulsion force, we examine the stability of the minimal model that couples the following dynamic variables: the two components of the polarization, $p_{x}$ and $p_{y}$, and the vorticity, $\Omega$. The vorticity is defined as $\Omega=\partial_{x}v_{y}-\partial_{y}v_{x}$ and alongside the steam function $\psi$ given by $v_{x}=\partial_{y}\psi$ and $v_{y}=-\partial_{x}\psi$ ensures the incompressibility condition. To isolate the essential mechanism underlying the instability, we neglect co-rotational terms that couple polarity dynamics to velocity gradients, as well as advective terms in the polarity and velocity equations. This simplification allows us to focus on the interplay between self-propulsion and flow alignment with the minimal ingredients required to trigger the instability. In the Appendix A, we show that including the full terms does not suppress the instability, underscoring the robustness of the underlying mechanism. Under these conditions, one can then write the Navier-Stokes equation (Eq. 2) in terms of vorticity
@@
This relation describes the evolution of the perturbation through a non-linear relation between the wave-number $q$ and the growth rate $\tilde{\omega}$ ( See Appendix C for the full expression). By differentiating Eq. (8) with respect to the wavenumber, we find that the growth rate reaches its maximum at the smallest wavenumber, $q=0$, similar to the hydrodynamic instabilities generated by dipolar flows in both polar [56] and nematic [57, 58] active systems. The fastest-growing mode corresponds to the wavenumber $q=\frac{2\pi}{L}$, specifically the longitudinal mode, with $n=1$ and $m=0$. In Fig. 3 we display the growth rate curves, for a constant self-propulsion coefficient and varying $\beta$. The critical self-propulsion $\alpha_{c}$ that separates the stable ($\tilde{\omega}<0$) from the unstable ($\tilde{\omega}>0$) states can be found by examining when $\tilde{\omega}$ crosses zero. This leads to the following relation
$\alpha_{c}=Kq^{2}\left(\frac{f}{\gamma}\frac{1}{\beta}+\beta\right),$ (9)
@@
Mechanism:- The relation that determines the systems stability, given by Eq. (9), reveals the three key components driving this instability: (i) the flow-aligning parameter $\beta$, (ii) the self-propulsion strength $\alpha$, and (iii) the Frank elastic constant $K$. Self-propulsion, $\alpha{\bf p}$, aligns the velocity with the polarization, and conversely, the coupling to the flow, $\beta\bf{v}$, aligns the polarization with the velocity (see schematic in Fig. 4a). When spatial variations in polarity occur, a competition arises between these aligning tendencies and the Frank elastic energy, which penalizes gradients in the polarization field (Fig. 4b). To illustrate this, consider an initially aligned velocity field with a polarity exhibiting a gradient along the $x$-axis (Fig. 4c). The self-propulsion and flow-alignment terms act to align velocity and polarity locally (Fig. 4d). However, this alignment frustrates the Frank elasticity, which favors a uniform polarization direction to minimize inhomogeneities. This competition creates a feedback loop: local misalignments lead to velocity changes that further distort the polarity field, amplifying fluctuations rather than suppressing them. Importantly, the instability arises because the flow alignment $\beta$ couples the polarity and velocity fields dynamically, and the absence of instantaneous force transmission allows these distortions to grow.
Discussion:- Our results demonstrate that a suspension of active self-propelled particles that are in contact with a substrate, the self-propulsive force alone is enough - in the absence of any dipolar active stress - to induce bend instabilities. This is noteworthy, since the hydrodynamics bend instability of a suspension of active particles is typically associated with active dipolar stresses.
@@
We note an important distinction between our results and those reported in Ref. [50]. Maitra et al. analyze fluctuations about a fully aligned polar state with a prescribed nonzero velocity, whereas here we perform the stability analysis about the isotropic, zero–velocity state. The latter is a natural reference configuration that does not consider perfect alignment. Furthermore, Ref. [50] eliminates velocity dynamics in the strict Stokes/adiabatic limit, while in our simulations we retain the unsteady term $\partial_{t}\vec{u}$ to track the finite relaxation of the flow and to decouple velocity–polarity modes. Crucially, this does not imply an inertia–driven mechanism: in all simulations $\mathrm{Re}<10^{-3}$. Retaining the unsteady term simply allows us to follow the temporal relaxation of the flow and to decouple velocity and polarity modes. Most importantly, our fully resolved simulations demonstrate that even when initialized in the aligned moving state considered by Ref. [50], the system destabilizes once fluctuations and nonlinearities are included.
@@
Our findings are potentially applicable to hydrodynamic instabilities and the early stages of collective motion in systems that share the essential physical ingredients captured in our model: polar order, self-propulsion, flow alignment, substrate friction, and hydrodynamic coupling within effectively incompressible films. This description encompasses a wide range of systems such as motility assays of polar filaments [13], migratory cells in monolayers [51, 10, 59], motile bacteria in two-dimensional setups [60, 61, 62], and active colloids [12, 38, 63] that move on substrates and tend to align with the local flow direction. In such systems, flow-alignment interactions are not merely a modeling assumption but arise naturally due to the presence of a physical substrate, which mediates frictional coupling and reinforces alignment between polarity and flow [64]. While prior work has often interpreted observed instabilities primarily through the lens of active stress–driven mechanisms, our analysis reveals a complementary and robust route to bend instabilities emerging from the interplay of self-propulsion and flow alignment. It is important to emphasize that this mechanism is distinct from the classic active nematic instability, which arises from gradients of dipolar active stress that generate destabilizing flows. These two routes to instability, one rooted in self-propulsion and flow alignment, the other in dipolar stress and flow gradient-alignment feedback, may coexist or compete, depending on system parameters. Distinguishing between them requires careful analysis of polarity dynamics and their relation to local flow versus strain rate fields. In this context, individual-based hydrodynamic models such as squirmers [65], which allow for independent tuning of self-propulsion and dipolar activity [66], offer a promising platform for systematically probing the interplay between these mechanisms. Likewise, direct measurements of active forces in experiments could help elucidate the relative contributions of polar and dipolar activity to the observed dynamics.
Finally, while the experimental confirmation of the polarity-driven instability mechanism has been limited by challenges in simultaneously quantifying polarity and flow fields, recent advances in imaging and analysis methods, specially in quantifying polarity of cells and bacteria [67, 68, 69] increasingly provide the tools to test these predictions directly.
@@ Appendix B Concentration effect on the linear stability
Moreover, we incorporate a term in the free energy that accounts for the coupling between the concentration and the splay deformation, $\mathcal{F}_{c}=-\int dA[w_{3}\frac{c-c_{0}}{c_{0}}\nabla\cdot p]$. Introducing concentration as a new dynamic variable and accounting for perturbations with respect to the mean concentration, $c=c_{0}+\delta c$, one can find the new Jacobian matrix,
@@
Appendix C Analytical growth rate relation of the main text
In this section we present the analytical expression of the growth rate that is discussed in the main text. The growth rate is obtained by solving the second-order algebraic equation in Eq. 8 of the main text, yielding the following expression:
$\color[rgb]{0,0,0}\tilde{\omega}_{\pm}=\frac{-\left(f+\rho\frac{K}{\gamma}q^{2}\right)\pm\sqrt{\left(f+\rho\frac{K}{\gamma}q^{2}\right)^{2}-4\rho\left[Kq^{2}\left(\frac{f}{\gamma}+\beta^{2}\right)-\beta\alpha\right]}}{2\rho}.$} (23)
In the main text, we focus on $\tilde{\omega}_{+}$ since it corresponds to the growth rate that can become positive.